OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[2\sqrt{50}-3\sqrt{8}\]

OpenStudy (anonymous):

what's the factor of 50?

OpenStudy (jamesj):

Ditto. If you understood your last two questions, you'd understand how to do this one.

OpenStudy (anonymous):

2 * sqrt25 * sqrt2 - thats the first bit

OpenStudy (lilg132):

oops \[10\sqrt{2}\] - \[6\sqrt{2}\]

OpenStudy (lilg132):

now this is very easy you just minus the two numbers outside of the \[\sqrt{}\]

OpenStudy (lilg132):

and you will get your answer im helping her learn here :p

OpenStudy (anonymous):

So my answer will be \[4\sqrt{2}\]? Or is that incorrect?

OpenStudy (lilg132):

well done

OpenStudy (anonymous):

Thanks lilg132 :))))))))

OpenStudy (lilg132):

see this is how you teach someone who isnt to keen on learning :p i should become a teacher

OpenStudy (lilg132):

your welcome :)

OpenStudy (anonymous):

Lol you should :P